OpenStudy (anonymous):

-2(z-4) - (3z-2) = -2 -(6z-2)

OpenStudy (anonymous):

-2(z-4)-(3z-2)=-2-(6z-2) -2z+8-3z+2=-2-6z+2 -5z+8=-6z -5z+6z=8 z=8

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

rain should the answer be - 8